### v3

# Fractional Sobolev logarithmic inequalities

Abstract.
We establish new Euclidean Sobolev logarithmic inequalities in the framework of fractional Sobolev spaces and their weighted version. Our approach relies on a interpolation inequality, which can be viewed as a fractional Caffarelli–Kohn–Nirenberg type inequality. We further relate the optimal constant in this interpolation inequality to a corresponding variational problem. These results extend classical Sobolev logarithmic inequalities to the nonlocal Euclidean framework and provide new tools for analysis in fractional Sobolev spaces.
Key words and phrases:
Sobolev logarithmic inequality, fractional Sobolev space
2020 Mathematics Subject Classification:
46E35, 26D15, 26A33
1. Introduction
This article focuses on proving a fractional version of the Euclidean $L^{p}$-Sobolev logarithmic inequality. In the local case, this inequality has been studied by many researchers. In particular, del Pino and Dolbeault [11] proved the Sobolev logarithmic inequality with the optimal constant. They showed that, for $1<p<d$, and for any $u\in W^{1,p}(\mathbb{R}^{d})$ with normalization $\int_{\mathbb{R}^{d}}|u(x)|^{p}\,dx=1$, we have
(1.1) $\int_{\mathbb{R}^{d}}|u(x)|^{p}\log|u(x)|\,dx\leq\frac{d}{p^{2}}\log\left[C_{p}\int_{\mathbb{R}^{d}}|\nabla u(x)|^{p}\,dx\right],$
where
(1.2) $C_{p}=\frac{p}{d}\left(\frac{p-1}{e}\right)^{p-1}\pi^{-\frac{p}{2}}\left[\frac{\Gamma\left(\frac{d}{2}+1\right)}{\Gamma\left(d\frac{p-1}{p}+1\right)}\right]^{\frac{p}{d}}$
is the best constant. Equality holds if and only if, for some $\sigma>0$ and $\tilde{x}\in\mathbb{R}^{d}$,
$u(x)=\pi^{\frac{d}{2}}\sigma^{-d\frac{(p-1)}{p}}\frac{\Gamma\left(\frac{d}{2}+1\right)}{\Gamma\left(d\frac{p-1}{p}+1\right)}e^{-\frac{1}{\sigma}|x-\tilde{x}|^{\frac{p}{p-1}}},\quad\forall\ x\in\mathbb{R}^{d}.$
This Sobolev logarithmic inequality is very important in the theory of partial differential equations. Since the Poincaré inequality in Sobolev spaces does not hold in $\mathbb{R}^{d}$, the above result can be seen as a natural analogue of the Sobolev logarithmic inequality in this case. For more work related to logarithmic Sobolev inequalities, we refer to [1, 3, 4, 7, 16, 21, 22].
Using the Sobolev inequality, one can also derive the Sobolev logarithmic inequality. However, in this case the constant obtained is not sharp. From the Sobolev inequality (see [4] and [11, Page No. 153]), we get
$\int_{\mathbb{R}^{d}}|u(x)|^{p}\log|u(x)|\,dx\leq\frac{d}{p^{2}}\log\left[D_{p}\int_{\mathbb{R}^{d}}|\nabla u(x)|^{p}\,dx\right],\quad\forall\ u\in W^{1,p}(\mathbb{R}^{d}),$
where $D_{p}$ is Talenti’s constant [20], given by
$D_{p}=\frac{1}{d}\left(\frac{p-1}{d-p}\right)^{p-1}\pi^{-\frac{p}{2}}\left(\frac{\Gamma\left(d\right)\Gamma\left(\frac{d}{2}+1\right)}{\Gamma\left(\frac{d}{p}\right)\Gamma\left(d\frac{p-1}{p}+1\right)}\right)^{\frac{p}{d}}.$
In this article, we establish a fractional version of the Sobolev logarithmic inequality of the type (1.1). For $p\geq 1$ and $s\in(0,1)$, the fractional Sobolev space $W^{s,p}(\mathbb{R}^{d})$ is defined as the completion of compactly supported smooth functions under the norm $(\|\cdot\|^{p}_{L^{p}(\mathbb{R}^{d})}+[\cdot]^{p}_{W^{s,p}(\mathbb{R}^{d})})^{\frac{1}{p}}$, where
$u]_{W^{s,p}(\mathbb{R}^{d})}:=\left(\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}\,dx\,dy\right)^{\frac{1}{p}}$
is called the Gagliardo seminorm. For more literature on fractional Sobolev spaces, we refer to the Hitchhiker’s guide [12].
The following theorem presents the main result of this article. We establish a fractional version of the Sobolev logarithmic inequality in the case $sp<d$, with an explicit constant. Furthermore, by letting $s\to 1$, the result reduces to the classical Sobolev logarithmic inequality. The precise statement is given below:
Theorem 1.1.
Let $p>1$ and $s\in(0,1)$ be such that $sp<d$. Then for every $u\in W^{s,p}(\mathbb{R}^{d})$ satisfying $\int_{\mathbb{R}^{d}}|u(x)|^{p}\,dx=1$, the following fractional Sobolev logarithmic inequality holds
(1.3) $\int_{\mathbb{R}^{d}}|u(x)|^{p}\log|u(x)|\,dx\leq\frac{d}{sp^{2}}\log\left(\mathcal{C}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}\,dx\,dy\right),$
where
(1.4) $\mathcal{C}=C(d,p)\left(\frac{s(1-s)}{(d-sp)^{p-1}}\right),$
with $C(d,p)>0$ depending only on $d$ and $p$.
Bourgain, Brezis, and Mironescu [5] established a fundamental link between fractional and classical Sobolev norms (see also [14, Theorem 3.19 and Corollary 3.20]). Specifically, for $p>1$ and $u\in L^{p}(\mathbb{R}^{d})$, there exists a constant $K=K(d,p)>0$ such that
(1.5) $\lim_{s\to 1}(1-s)\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}\,dx\,dy=K\int_{\mathbb{R}^{d}}|\nabla u(x)|^{p}\,dx.$
Using this relation, and letting $s\to 1$ in inequality (1.3) of Theorem 1.1, we recover the classical Sobolev logarithmic inequality. In particular, we obtain
$\int_{\mathbb{R}^{d}}|u(x)|^{p}\log|u(x)|\,dx\leq\frac{d}{p^{2}}\log\left(C(d,p)\frac{K}{(d-p)^{p-1}}\int_{\mathbb{R}^{d}}|\nabla u(x)|^{p}\,dx\right).$
del Pino and Dolbeault [11] derived the same inequality with the sharp constant $C_{p}$ given in (1.2). Consequently, we deduce the bound
$C_{p}\leq C(d,p)\frac{K}{(d-p)^{p-1}}.$
A version of the above theorem on stratified Lie groups was proved by Ghosh, Kumar, and Ruzhansky in [15, Section $7$]. They established subelliptic fractional logarithmic Sobolev inequalities on stratified Lie groups with explicit constants depending on an additional parameter lying between $p$ and the fractional critical Sobolev exponent, under the condition $sp<d$. In contrast, Theorem 1.1 provides a constant $\mathcal{C}$ in Euclidean space that is independent of any extra parameter and also recovers the classical Sobolev logarithmic inequality.
Lieb and Loss, in [17], proved another version of the Sobolev logarithmic inequality. It says that for any $c>0$, we have
(1.6) $\int_{\mathbb{R}^{d}}|u(x)|^{2}\log\left(\frac{|u(x)|^{2}}{\|u\|^{2}_{L^{2}(\mathbb{R}^{d})}}\right)\,dx+d(1+\log c)\|u\|^{2}_{L^{2}(\mathbb{R}^{d})}\leq\frac{c^{2}}{\pi}\int_{\mathbb{R}^{d}}|\nabla u(x)|^{2}\,dx.$
This inequality becomes an equality only when $u$ is a constant multiple of the Gaussian function $\exp\left(-\frac{\pi|x|^{2}}{2c^{2}}\right)$. Later, Chatzakou and Ruzhansky, in [8], proved a fractional version of this inequality. They showed that if $u\in W^{s,2}(\mathbb{R}^{d})$ with $0<s<\frac{n}{2}$ and $c>0$, then
$\int_{\mathbb{R}^{d}}|u(x)|^{2}\log\left(\frac{|u(x)|^{2}}{\|u\|^{2}_{L^{2}(\mathbb{R}^{d})}}\right)\,dx+\frac{d}{s}(1+\log c)\|u\|^{2}_{L^{2}(\mathbb{R}^{d})}\leq\frac{dec^{2}}{2s}C_{d,s}\|(-\Delta u)^{\frac{s}{2}}u\|^{2}_{L^{2}(\mathbb{R}^{d})},$
where
$C_{d,s}=\frac{\Gamma\left(\frac{d-2s}{2}\right)}{2^{2s}\pi^{s}\Gamma\left(\frac{d+2s}{2}\right)}\left(\frac{\Gamma\left(d\right)}{\Gamma\left(\frac{d}{2}\right)}\right)^{\frac{2s}{d}}$
and
$\|(-\Delta u)^{\frac{s}{2}}u\|^{2}_{L^{2}(\mathbb{R}^{d})}=\int_{\mathbb{R}^{d}}|\widehat{u}(\xi)|^{2}(4\pi^{2}|\xi|^{2})^{s}d\xi.$
Here, $\widehat{u}$ denotes the Fourier transform of $u$.
In the next theorem, we extend the fractional Sobolev logarithmic inequality established in [8] to the case $sp<d$, with an explicit constant. Moreover, by passing to the limit $s\to 1$, we recover the another form of classical Sobolev logarithmic inequality.
Theorem 1.2.
Let $p>1$ and $s\in(0,1)$ be such that $sp<d$. Then for any $u\in W^{s,p}(\mathbb{R}^{d})$ and any real number $c>0$, the following inequality holds
$\int_{\mathbb{R}^{d}}|u(x)|^{p}\log\left(\frac{|u(x)|^{p}}{\|u\|^{p}_{L^{p}(\mathbb{R}^{d})}}\right)\,dx\,+\,$ $\frac{d}{s}(1+\log c)\|u\|^{p}_{L^{p}(\mathbb{R}^{d})}$
(1.7) $\leq\frac{de^{p-1}c^{p}\,\mathcal{C}}{sp}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}\,dx\,dy,$
where $\mathcal{C}$ is as defined in (2.2).
By applying the relation (1.5) in the above theorem, we recover the local case. In particular, taking the limit $s\to 1$ in the above theorem, we obtain the following another form of the Sobolev logarithmic inequality in the local setting:
$\int_{\mathbb{R}^{d}}|u(x)|^{p}\log\left(\frac{|u(x)|^{p}}{\|u\|^{p}_{L^{p}(\mathbb{R}^{d})}}\right)\,dx\,+\,$ $d(1+\log c)\|u\|^{p}_{L^{p}(\mathbb{R}^{d})}$
(1.8) $\leq\frac{de^{p-1}c^{p}\,K}{p}\left(\frac{C(d,p)}{(d-p)^{p-1}}\right)\int_{\mathbb{R}^{d}}|\nabla u(x)|^{p}\,dx.$
The next theorem establishes a fractional Sobolev logarithmic inequality in the weighted fractional Sobolev space. We consider the space $W^{s,p,\alpha}(\mathbb{R}^{d})$ as the closure of $C^{1}_{c}(\mathbb{R}^{d})$ with respect to the weighted norm $\|\cdot\|_{W^{s,p,\alpha}(\mathbb{R}^{d})}$ (see (3.1) for the precise definition). Here $\alpha,\alpha_{1},\alpha_{2}\in\mathbb{R}$ are parameters with $\alpha=\alpha_{1}+\alpha_{2}$ together with certain admissible bounds. The main result is stated as follows:
Theorem 1.3.
Let $p>1$, $s\in(0,1)$, and $\alpha_{1},\alpha_{2}\in\mathbb{R}$ be such that $\alpha=\alpha_{1}+\alpha_{2}$, $\alpha_{1}p,\,\alpha_{2}p\in(-d,sp)$, $0\leq\alpha p<sp$, and $sp-\alpha p<d$. Then for every $u\in W^{s,p,\alpha}(\mathbb{R}^{d})$ satisfying $\int_{\mathbb{R}^{d}}|u(x)|^{p}\,dx=1$, the following weighted fractional Sobolev logarithmic inequality holds
(1.9) $\int_{\mathbb{R}^{d}}|u(x)|^{p}\log|u(x)|\,dx\leq\frac{d}{(s-\alpha)p^{2}}\log\left(\mathcal{C}_{\alpha}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}|x|^{\alpha_{1}p}|y|^{\alpha_{2}p}\,dx\,dy\right),$
where $\mathcal{C}_{\alpha}=\mathcal{C}_{\alpha}(d,p,s,\alpha_{1},\alpha_{2})>0$.
The key step in proving the above three theorems is to establish an interpolation inequality, or in other words, a fractional Caffarelli–Kohn–Nirenberg type inequality. In Section 2 and Section 3, we will derive such an interpolation inequality with a constant. In the next theorem, we state an interpolation inequality in the fractional Sobolev space with an optimal constant $\mathscr{L}$. This optimal constant $\mathscr{L}$ is connected to the infimum of a variational problem, which we will discuss in Section 4. For the literature on Caffarelli–Kohn–Nirenberg type inequalities, we refer to [2, 6, 9, 19]. For $s\in(0,1)$, $p>1$, and $q>p$, we define the space $\mathcal{D}^{p,q}_{s}$ as
(1.10) $\mathcal{D}^{p,q}_{s}=(W^{s,p}(\mathbb{R}^{d})\cap L^{q}(\mathbb{R}^{d}))\setminus\{0\}.$
The next interpolation theorem is stated as follows.
Theorem 1.4.
Let $p>1$ and $s\in(0,1)$ be such that $sp<d$, $p<q<\frac{p(d-s)}{d-sp}$, and define $r=p\,\frac{q-1}{p-1}$. Then for every $u\in\mathcal{D}^{p,q}_{s}$, the following interpolation inequality holds with an optimal constant $\mathscr{L}$,
(1.11) $\|u\|_{L^{r}(\mathbb{R}^{d})}\leq\mathscr{L}\,[u]^{a}_{W^{s,p}(\mathbb{R}^{d})}\,\|u\|^{1-a}_{L^{q}(\mathbb{R}^{d})},$
where
(1.12) $a=\frac{d(q-p)}{(q-1)\bigl(dp-(d-sp)q\bigr)},$
and $\delta:=dp-(d-sp)q>0$.
This article is organized as follows. In Section 2, we prove Theorems 1.1 and 1.2, and establish an interpolation inequality that plays a central role in these proofs. Section 3 is devoted to extending logarithmic Sobolev inequalities to weighted fractional Sobolev spaces, where we prove Theorem 1.3. Finally, in Section 4, we relate the optimal constant in the interpolation inequality to the infimum of a corresponding variational problem.
2. Proof of fractional Sobolev logarithmic inequalities
In this section, we prove Theorem 1.1 and Theorem 1.2 by establishing the fractional Sobolev logarithmic inequalities for the case $sp<d$. The main tool is the fractional Sobolev inequality with an explicit constant, given by Maz’ya and Shaposhnikova in [18, Theorem 1]. The constant $\mathcal{C}$ in Theorem 1.1 is the same as the one in [18, Theorem 1]. Using this fractional Sobolev inequality, we derive an interpolation inequality with the constant $\mathcal{C}^{\frac{a}{p}}$. However, this constant $\mathcal{C}^{\frac{a}{p}}$ is not sharp. In Section 4, we will show the relation between the best constant in the interpolation inequality and a corresponding variational problem.
Next, we recall the fractional Sobolev inequality in the case $sp<d$, where $p^{*}_{s}:=\frac{dp}{d-sp}$. This result is also due to Maz’ya and Shaposhnikova [18, Theorem 1]. For the case $p=2$, the optimal constant in fractional Sobolev inequality was established by Cotsiolis and Tavoularis in [10].
Lemma 2.1.
Let $d\geq 1$, $p\geq 1$ and $s\in(0,1)$ such that $sp<d$. Then for any $u\in W^{s,p}(\mathbb{R}^{d})$, there holds
(2.1) $\left(\int_{\mathbb{R}^{d}}|u(x)|^{p^{*}_{s}}\,dx\right)^{\frac{1}{p^{*}_{s}}}\leq\mathcal{C}^{\frac{1}{p}}\left(\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}\,dx\,dy\right)^{\frac{1}{p}},$
where $p^{*}_{s}=\frac{dp}{d-sp}$, and
(2.2) $\mathcal{C}=C(d,p)\left(\frac{s(1-s)}{(d-sp)^{p-1}}\right),$
with $C(d,p)>0$ depending only on $d$ and $p$.
Now we derive an interpolation inequality in the fractional Sobolev space. For this, we make use of the above lemma. This result is an important step towards proving Theorem 1.1 and Theorem 1.2.
Theorem 2.2.
Let $p>1$ and $s\in(0,1)$ be such that $sp<d$, $p<q<\tfrac{p(d-s)}{d-sp}$, and $r=p\tfrac{q-1}{p-1}$. Then for all $u\in\mathcal{D}^{p,q}_{s}$,
(2.3) $\|u\|_{L^{r}(\mathbb{R}^{d})}\leq\mathcal{C}^{\frac{a}{p}}[u]^{a}_{W^{s,p}(\mathbb{R}^{d})}\|u\|_{L^{q}(\mathbb{R}^{d})}^{1-a},$
where $\mathcal{C}$ is as defined in (2.2), and
(2.4) $a=\frac{d(q-p)}{(q-1)(dp-(d-sp)q)},$
with $\delta=dp-(d-sp)q>0$.
Proof.
It is easy to see that $q<r<p^{*}_{s}$. Next, we determine the interpolation parameter $a\in(0,1)$ by requiring
$\frac{1}{r}=\frac{a}{p^{*}_{s}}+\frac{1-a}{q}.$
Solving for $a$, we obtain
$a=\frac{\tfrac{1}{r}-\tfrac{1}{q}}{\tfrac{1}{p^{*}_{s}}-\tfrac{1}{q}}=\frac{\dfrac{q-r}{rq}}{\dfrac{q-p^{*}_{s}}{p^{*}_{s}q}}=\frac{(q-r)p^{*}_{s}}{(q-p^{*}_{s})r}.$
Since
$q-r=q-\frac{p(q-1)}{p-1}=\frac{q(p-1)-p(q-1)}{p-1}=\frac{p-q}{p-1},$
and
$q-p^{*}_{s}=q-\frac{dp}{d-sp}=\frac{q(d-sp)-dp}{d-sp}=-\frac{dp-(d-sp)q}{d-sp},$
we deduce
$a=\frac{p^{*}_{s}\tfrac{p-q}{p-1}}{r\big(-\tfrac{dp-(d-sp)q}{d-sp}\big)}=\frac{p^{*}_{s}(q-p)(d-sp)}{(p-1)r\,(dp-(d-sp)q))}.$
Since $p^{*}_{s}(d-sp)=dp$ and (p1)r=p(q1)(p-1)r=p(q-1), this reduces to
$a=\frac{d(q-p)}{(q-1)\,(dp-(d-sp)q)}.$
Now we estimate $\|u\|_{L^{r}(\mathbb{R}^{d})}$. We write
$\|u\|^{r}_{L^{r}(\mathbb{R}^{d})}=\int_{\mathbb{R}^{d}}|u(x)|^{r}\,dx=\int_{\mathbb{R}^{d}}|u(x)|^{ar}\,|u(x)|^{(1-a)r}\,dx.$
Applying Hölder’s inequality with exponents
$\nu=\frac{p^{*}_{s}}{ar},\qquad\xi=\frac{q}{(1-a)r},$
we obtain
$\int_{\mathbb{R}^{d}}|u(x)|^{ar}|u(x)|^{(1-a)r}\,dx\leq\|u\|_{L^{p^{*}_{s}}}^{ar}\,\|u\|_{L^{q}}^{(1-a)r}.$
Thus,
$\|u\|_{L^{r}(\mathbb{R}^{d})}\leq\|u\|_{L^{p^{*}_{s}}(\mathbb{R}^{d})}^{a}\,\|u\|_{L^{q}(\mathbb{R}^{d})}^{1-a}.$
Finally, using the fractional Sobolev inequality (see Lemma 2.1), we conclude
$\|u\|_{L^{r}(\mathbb{R}^{d})}\leq\mathcal{C}^{\frac{a}{p}}[u]^{a}_{W^{s,p}(\mathbb{R}^{d})}\|u\|_{L^{q}(\mathbb{R}^{d})}^{1-a}.$
This proves the theorem. ∎
The following lemma was used by del Pino and Dolbeault in [11] to prove the Sobolev logarithmic inequality in the local case. They combined an interpolation inequality with this lemma to obtain their result. We also use the same lemma to establish the logarithmic Sobolev inequality in the non-local case. For completeness, we give the proof of the following lemma.
Lemma 2.3.
Let $1\leq q<r<\infty$. Then for any $u\in L^{q}(\mathbb{R}^{d})\cap L^{r}(\mathbb{R}^{d})$, we have
(2.5) $\int_{\mathbb{R}^{d}}|u(x)|^{q}\log\left(\frac{|u(x)|}{\|u\|_{L^{q}(\mathbb{R}^{d})}}\right)\,dx\leq\frac{r}{r-q}\|u\|^{q}_{L^{q}(\mathbb{R}^{d})}\log\left(\frac{\|u\|_{L^{r}(\mathbb{R}^{d})}}{\|u\|_{L^{q}(\mathbb{R}^{d})}}\right).$
Proof.
By applying Hölder’s inequality with $\alpha(s)=\frac{q}{s}\left(\frac{r-s}{r-q}\right)$, $q\leq s\leq r$, we get
$\|u\|_{L^{s}(\mathbb{R}^{d})}\leq\|u\|^{\alpha(s)}_{L^{q}(\mathbb{R}^{d})}\|u\|^{1-\alpha(s)}_{L^{r}(\mathbb{R}^{d})}.$
Taking logarithm on both sides,
(2.6) $\log\|u\|_{L^{s}(\mathbb{R}^{d})}\leq\alpha(s)\log\|u\|_{L^{q}(\mathbb{R}^{d})}+(1-\alpha(s))\log\|u\|_{L^{r}(\mathbb{R}^{d})}.$
Now, define $\Phi(s):=\log\|u\|_{L^{s}(\mathbb{R}^{d})}=\frac{1}{s}\log\left(\int_{\mathbb{R}^{d}}|u(x)|^{s}\,dx\right)$. Then, we have
(2.7) $s\Phi^{\prime}(s)=\frac{1}{\|u\|^{s}_{L^{s}(\mathbb{R}^{d})}}\int_{\mathbb{R}^{d}}|u(x)|^{s}\log\left(\frac{|u(x)|}{\|u\|_{L^{s}(\mathbb{R}^{d})}}\right)\,dx.$
Also, set $\theta(s)=\alpha(s)\log\|u\|_{L^{q}(\mathbb{R}^{d})}+(1-\alpha(s))\log\|u\|_{L^{r}(\mathbb{R}^{d})}$. Then,
$\theta^{\prime}(s)$ $=\alpha^{\prime}(s)\left(\log\|u\|_{L^{q}(\mathbb{R}^{d})}-\log\|u\|_{L^{r}(\mathbb{R}^{d})}\right)$
(2.8) $=\left(\frac{qr}{s^{2}(r-q)}\right)\log\left(\frac{\|u\|_{L^{r}(\mathbb{R}^{d})}}{\|u\|_{L^{q}(\mathbb{R}^{d})}}\right).$
From inequality (2.6), we have $\Phi(s)\leq\theta(s)$ for $q\leq s\leq r$, and equality when $s=q$. So, $\Phi^{\prime}(q)\leq\theta^{\prime}(q)$. Combining (2.7) and (2) at $s=q$, we obtain
$\int_{\mathbb{R}^{d}}|u(x)|^{q}\log\left(\frac{|u(x)|}{\|u\|_{L^{q}(\mathbb{R}^{d})}}\right)\,dx\leq\frac{r}{r-q}\|u\|^{q}_{L^{q}(\mathbb{R}^{d})}\log\left(\frac{\|u\|_{L^{r}(\mathbb{R}^{d})}}{\|u\|_{L^{q}(\mathbb{R}^{d})}}\right).$
Hence, the lemma is proved. ∎
Now we prove Theorem 1.1 by combining Theorem 2.2 with Lemma 2.3. Note that the constant $\mathcal{C}$ in our fractional Sobolev logarithmic inequality is not optimal.
Proof of Theorem 1.1.
Let us assume that $sp<d,~p<q<\frac{p(d-s)}{d-sp}$, and $r=p\frac{q-1}{p-1}$. From Theorem 2.2, we have
$\|u\|_{L^{r}(\mathbb{R}^{d})}\leq\mathcal{C}^{\frac{a}{p}}[u]^{a}_{W^{s,p}(\mathbb{R}^{d})}\|u\|_{L^{q}(\mathbb{R}^{d})}^{1-a},\quad\forall\ u\in\mathcal{D}^{p,q}_{s},$
where
$a=\frac{d(q-p)}{(q-1)(dp-(d-sp)q)}.$
The above inequality can be equivalently written as
$\frac{\|u\|_{L^{r}(\mathbb{R}^{d})}}{\|u\|_{L^{q}(\mathbb{R}^{d})}}\leq\mathcal{C}^{\frac{a}{p}}\left(\frac{[u]_{W^{s,p}(\mathbb{R}^{d})}}{\|u\|_{L^{q}(\mathbb{R}^{d})}}\right)^{a}.$
Taking logarithms on both sides yields
(2.9) $\frac{1}{a}\log\left(\frac{\|u\|_{L^{r}(\mathbb{R}^{d})}}{\|u\|_{L^{q}(\mathbb{R}^{d})}}\right)-\log\left(\mathcal{C}^{\frac{1}{p}}\right)\leq\log\left(\frac{[u]_{W^{s,p}(\mathbb{R}^{d})}}{\|u\|_{L^{q}(\mathbb{R}^{d})}}\right).$
From Lemma 2.3, we further have
$\left(\frac{r-q}{r}\right)\frac{1}{a}\int_{\mathbb{R}^{d}}\frac{|u(x)|^{q}}{\|u\|^{q}_{L^{q}(\mathbb{R}^{d})}}\log\left(\frac{|u(x)|}{\|u\|_{L^{q}(\mathbb{R}^{d})}}\right)\,dx\leq\frac{1}{a}\log\left(\frac{\|u\|_{L^{r}(\mathbb{R}^{d})}}{\|u\|_{L^{q}(\mathbb{R}^{d})}}\right).$
Combining the above two inequalities and letting $q\to p$, with the values of $r$ and $a$ given in Theorem 2.2, we obtain
(2.10) $\frac{sp}{d}\int_{\mathbb{R}^{d}}\frac{|u(x)|^{p}}{\|u\|^{p}_{L^{p}(\mathbb{R}^{d})}}\log\left(\frac{|u(x)|}{\|u\|_{L^{p}(\mathbb{R}^{d})}}\right)\,dx-\log\left(\mathcal{C}^{\frac{1}{p}}\right)\leq\log\left(\frac{[u]_{W^{s,p}(\mathbb{R}^{d})}}{\|u\|_{L^{p}(\mathbb{R}^{d})}}\right).$
Finally, imposing the normalization $\int_{\mathbb{R}^{d}}|u(x)|^{p}\,dx=1$, we deduce
[…]
This completes the proof. ∎
Proof of Theorem 1.2.
We begin with the assumption of Theorem 1.2. From inequality (2.10) in the proof of Theorem 1.1, we know that
$\frac{sp}{d}\int_{\mathbb{R}^{d}}\frac{|u(x)|^{p}}{\|u\|^{p}_{L^{p}(\mathbb{R}^{d})}}\log\left(\frac{|u(x)|}{\|u\|_{L^{p}(\mathbb{R}^{d})}}\right)\,dx-\log\left(\mathcal{C}^{\frac{1}{p}}\right)\leq\log\left(\frac{[u]_{W^{s,p}(\mathbb{R}^{d})}}{\|u\|_{L^{p}(\mathbb{R}^{d})}}\right).$
This inequality can be rewritten as
[…]
It is easy to check that for any $b>0$ and $x>0$, the following inequality holds true
$\log x\leq bx-\log b-1.$
[…]
$\hskip 17.07164pt-(\log(e^{p-1}c^{p})+1)\Big)$
$=e^{p-1}c^{p}\,\mathcal{C}[u]^{p}_{W^{s,p}(\mathbb{R}^{d})}-p(\log c+1)\|u\|^{p}_{L^{p}(\mathbb{R}^{d})}.$
Therefore, we arrive at
$\int_{\mathbb{R}^{d}}|u(x)|^{p}\log\left(\frac{|u(x)|^{p}}{\|u\|^{p}_{L^{p}(\mathbb{R}^{d})}}\right)\,dx+\frac{d}{s}(1+\log c)\|u\|^{p}_{L^{p}(\mathbb{R}^{d})}\leq\frac{de^{p-1}c^{p}\,\mathcal{C}}{sp}[u]^{p}_{W^{s,p}(\mathbb{R}^{d})}.$
This finishes the proof of the theorem. ∎
3. Proof of Weighted fractional Sobolev logarithmic inequality
In this section, we extend the fractional Sobolev logarithmic inequalities to the weighted fractional Sobolev space and establish Theorem 1.3. We begin by defining the weighted fractional Sobolev space. For $p>1$, $s\in(0,1)$, and $\alpha,\alpha_{1},\alpha_{2}\in\mathbb{R}$ with $\alpha=\alpha_{1}+\alpha_{2}$, the weighted Gagliardo seminorm is given by
$u]_{W^{s,p,\alpha}(\mathbb{R}^{d})}:=\left(\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}|x|^{\alpha_{1}p}|y|^{\alpha_{2}p}\,dx\,dy\right)^{\frac{1}{p}}\leq\infty,\quad\forall\ u\in L^{p}(\mathbb{R}^{d}).$
The norm of this space is defined as
(3.1) $\|u\|_{W^{s,p,\alpha}(\mathbb{R}^{d})}=\left(\|u\|^{p}_{L^{p}(\mathbb{R}^{d})}+[u]^{p}_{W^{s,p,\alpha}(\mathbb{R}^{d})}\right)^{\frac{1}{p}}.$
Dipierro and Valdinoci [13] proved that if $\alpha_{1}p,\,\alpha_{2}p\in(-d,sp)$ and $\alpha_{1}p+\alpha_{2}p>-d$, then for every $u\in C^{1}_{c}(\mathbb{R}^{d})$, the seminorm $u]_{W^{s,p,\alpha}(\mathbb{R}^{d})}<\infty$. We then define the space $W^{s,p,\alpha}(\mathbb{R}^{d})$ as the closure of $C^{1}_{c}(\mathbb{R}^{d})$ with respect to the norm $\|\cdot\|_{W^{s,p,\alpha}(\mathbb{R}^{d})}$.
Nguyen and Squassina in [19] studied Caffarelli–Kohn–Nirenberg type inequalities in fractional Sobolev spaces and proved the following theorem. Let $d\geq 1$, $p>1$, $q\geq 1$, $\tau>0$, $a\in(0,1]$, and $\alpha,\beta,\gamma\in\mathbb{R}$ satisfy
$\frac{1}{\tau}+\frac{\gamma}{d}=a\left(\frac{1}{p}+\frac{\alpha-s}{d}\right)+(1-a)\left(\frac{1}{q}+\frac{\beta}{d}\right)\ .$
If $a>0$, assume also that, with $\gamma=a\sigma+(1-a)\beta$,
$0\leq\alpha-\sigma$
and
$\alpha-\sigma\leq s\hskip 14.22636pt\text{if}\hskip 14.22636pt\frac{1}{\tau}+\frac{\gamma}{d}=\frac{1}{p}+\frac{\alpha-s}{d}.$
If $\frac{1}{\tau}+\frac{\gamma}{d}>0$, then we have
(3.2) $\||x|^{\gamma}u\|_{L^{\tau}(\mathbb{R}^{d})}\leq C[u]^{a}_{W^{s,p,\alpha}(\mathbb{R}^{d})}\||x|^{\beta}u\|^{(1-a)}_{L^{q}(\mathbb{R}^{d})},\hskip 8.5359pt\forall\ u\in C^{1}_{c}(\mathbb{R}^{d}).$
Now, if we take $\gamma=0$ and $a=1$, the fractional Caffarelli–Kohn–Nirenberg inequality (3.2) reduces to
(3.3) $\left(\int_{\mathbb{R}^{d}}|u(x)|^{\tau}\,dx\right)^{\!\frac{1}{\tau}}\leq C\left(\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}|x|^{\alpha_{1}p}|y|^{\alpha_{2}p}\,dx\,dy\right)^{\!\frac{1}{p}},$
under the conditions
$0\leq\alpha<s,\qquad sp-\alpha p<d,\qquad\text{and}\qquad\tau=\frac{dp}{d-sp+\alpha p}.$
The inequality (3.3) is nothing but the fractional Sobolev inequality in weighted spaces, with the critical exponent
$p^{*}_{s,\alpha}:=\tau=\frac{dp}{d-sp+\alpha p},\qquad\text{if}\quad sp-\alpha p<d.$
We can therefore state the following fundamental result.
Lemma 3.1 (Weighted fractional Sobolev inequality).
Let $p>1$, $s\in(0,1)$, and $\alpha_{1},\alpha_{2}\in\mathbb{R}$ be such that $\alpha=\alpha_{1}+\alpha_{2}$, $\alpha_{1}p$, $\alpha_{2}p\in(-d,sp)$, $0\leq\alpha p<sp$, $sp-\alpha p<d$ with $p^{*}_{s,\alpha}=dp/(d-sp+\alpha p)$. Then there exists a constant $\mathcal{C}_{\alpha}=\mathcal{C}_{\alpha}(d,p,s,\alpha_{1},\alpha_{2})>0$ such that
$\left(\int_{\mathbb{R}^{d}}|u(x)|^{p^{*}_{s,\alpha}}\,dx\right)^{\!\frac{1}{p^{*}_{s,\alpha}}}\leq\mathcal{C}^{\frac{1}{p}}_{\alpha}\Big(\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}|x|^{\alpha_{1}p}|y|^{\alpha_{2}p}$ $\,dx\,dy\Big)^{\!\frac{1}{p}},$
(3.4) $\forall\ u\in W^{s,p,\alpha}(\mathbb{R}^{d}).$
We now derive a weighted version of the interpolation inequality, as formulated in Theorem 2.2. For $s\in(0,1)$, $p>1$, $\alpha=\alpha_{1}+\alpha_{2}$, $\alpha_{1}p$, $\alpha_{2}p\in(-d,sp)$, $\alpha_{1}p+\alpha_{2}p>-d$, $sp-\alpha p<d$, and $q>p$, we define the space $\mathcal{D}^{p,q}_{s,\alpha}$ as
(3.5) $\mathcal{D}^{p,q}_{s,\alpha}=(W^{s,p,\alpha}(\mathbb{R}^{d})\cap L^{q}(\mathbb{R}^{d}))\setminus\{0\}.$
Following the strategy used in the proof of Theorem 2.2, we obtain the following result:
Theorem 3.2.
Let $p>1$, $s\in(0,1)$, and $\alpha_{1},\alpha_{2}\in\mathbb{R}$ be such that $\alpha=\alpha_{1}+\alpha_{2}$, $\alpha_{1}p$, $\alpha_{2}p\in(-d,sp)$, $0\leq\alpha p<sp$, and $sp-\alpha p<d$. Let $q$ be such that $p<q<\frac{p(d-s+\alpha)}{d-sp+\alpha p}$, and set $r=p\frac{q-1}{p-1}$. Then there exists a constant $\mathcal{C}_{\alpha}=\mathcal{C}_{\alpha}(d,p,s,\alpha_{1},\alpha_{2})>0$ such that, for all $u\in\mathcal{D}^{p,q}_{s,\alpha}$,
(3.6) $\|u\|_{L^{r}(\mathbb{R}^{d})}\leq\mathcal{C}_{\alpha}^{\frac{a}{p}}[u]_{W^{s,p,\alpha}(\mathbb{R}^{d})}^{a}\|u\|_{L^{q}(\mathbb{R}^{d})}^{1-a},$
where $\mathcal{C}_{\alpha}$ is as defined in Lemma 3.1, and
(3.7) $a=\frac{d(q-p)}{(q-1)\big(dp-(d-sp+\alpha p)q\big)},$
with $\delta_{\alpha}:=dp-(d-sp+\alpha p)q>0$.
Proof of Theorem 1.3.
Following the strategy of the proof of Theorem 1.1, and passing to the limit $q\to p$, we obtain that under the normalization $\int_{\mathbb{R}^{d}}|u(x)|^{p}\,dx=1$, the following inequality holds:
$\int_{\mathbb{R}^{d}}|u(x)|^{p}\log|u(x)|\,dx\leq\frac{d}{(s-\alpha)p^{2}}\log\left(\mathcal{C}_{\alpha}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}|x|^{\alpha_{1}p}|y|^{\alpha_{2}p}\,dx\,dy\right).$
This completes the proof. ∎
Remark 1.
By adapting the proof of Theorem 1.2, we also obtain the following inequality. For any $c>0$ and any $u\in W^{s,p,\alpha}(\mathbb{R}^{d})$,
$\int_{\mathbb{R}^{d}}|u(x)|^{p}\log\left(\frac{|u(x)|^{p}}{\|u\|^{p}_{L^{p}(\mathbb{R}^{d})}}\right)dx$ $+\frac{d}{s-\alpha}(1+\log c)\,\|u\|^{p}_{L^{p}(\mathbb{R}^{d})}$
(3.8) $\leq\frac{d\,e^{p-1}c^{p}\,\mathcal{C}_{\alpha}}{(s-\alpha)p}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}|x|^{\alpha_{1}p}|y|^{\alpha_{2}p}\,dx\,dy.$
4. Proof of Theorem 1.4
In this section, we prove Theorem 1.4 with the optimal constant $\mathscr{L}$. We begin by setting up a variational problem, then connect it to the interpolation inequality, and finally show that the infimum value of this variational problem gives the best constant in the inequality.
The next lemma explains the connection between the variational problem and the interpolation inequality with the optimal constant.
Lemma 4.1.
For any fixed a,b,M,N>0a,~b,~M,~N>0, define $f(s)=s^{a}M+s^{-b}N$, for all $s>0$. Then the minimum value of $f(s)$ is attained at $s=\left(\frac{bN}{aM}\right)^{\frac{1}{a+b}}$, and this minimum value is
(4.1) $\inf_{s>0}f(s)=f\left(\left(\frac{bN}{aM}\right)^{\frac{1}{a+b}}\right)=\frac{a+b}{a}\left(\frac{b}{a}\right)^{\frac{-b}{a+b}}M^{\frac{b}{a+b}}N^{\frac{a}{a+b}}.$
Proof.
We first observe that
$\lim_{s\to 0^{+}}f(s)=\infty\quad\text{and}\quad\lim_{s\to\infty}f(s)=\infty.$
Hence the function $f(s)$ must attain its minimum at some point $s>0$. Differentiating, we get
$f^{\prime}(s)=aMs^{a-1}-bNs^{-b-1}.$
Setting $f^{\prime}(s)=0$ gives
$aMs^{a-1}=bNs^{-b-1}\quad\Longrightarrow\quad s^{a+b}=\frac{bN}{aM}.$
Thus the only critical point is
$s_{*}=\left(\frac{bN}{aM}\right)^{\tfrac{1}{a+b}}.$
To check that this corresponds to a minimum, note that for $s\to 0^{+}$ or $s\to\infty$ we have $f(s)\to\infty$, so the finite critical point $s_{*}$ must indeed give the global minimum. Finally, substituting $s_{*}$ into $f(s)$ we obtain
$f(s_{*})=M\left(\frac{bN}{aM}\right)^{\tfrac{a}{a+b}}+N\left(\frac{bN}{aM}\right)^{-\tfrac{b}{a+b}}=\frac{a+b}{a}\left(\frac{b}{a}\right)^{-\tfrac{b}{a+b}}M^{\tfrac{b}{a+b}}N^{\tfrac{a}{a+b}}.$
This proves the claim. ∎
Proof of Theorem 1.4.
Under the assumption of Theorem 1.4, define the functional
(4.2) $I(u):=\frac{1}{p}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}\,dx\,dy+\frac{1}{q}\int_{\mathbb{R}^{d}}|u(x)|^{q}\,dx.$
Now, let $K>0$ be a constant, which we fix later. Conside the set
$\Lambda:=\left\{u\in\mathcal{D}^{p,q}_{s}:\int_{\mathbb{R}^{d}}|u(x)|^{r}\,dx=K\right\}.$
From Theorem 2.2, we immediately get
(4.3) $\eta:=\inf_{u\in\Lambda}I(u)>0.$
Next, define the scaling transform for parameters $\alpha:=\frac{d(p-1)}{p(q-1)}>0$ and $\lambda>0$,
$T_{\lambda}(u)=u_{\lambda}(x)=\lambda^{\alpha}u(\lambda x).$
By change of variables, we find
$\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u_{\lambda}(x)-u_{\lambda}(y)|^{p}}{|x-y|^{d+sp}}\,dx\,dy=\lambda^{\alpha p-(d-sp)}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}\,dx\,dy,$
and
$\int_{\mathbb{R}^{d}}|u_{\lambda}(x)|^{q}\,dx=\lambda^{-(d-\alpha q)}\int_{\mathbb{R}^{d}}|u(x)|^{q}\,dx.$
By using the definition of $\alpha$, we have
$\alpha p-(d-sp)=\frac{p(d-s)-q(d-sp)}{q-1}>0\hskip 8.53581pt\text{iff}\hskip 8.53581ptq<\frac{p(d-s)}{d-sp},$
and
$d-\alpha q=\frac{d(q-p)}{p(q-1)}>0\hskip 8.53581pt\text{as}\hskip 8.53581ptq>p.$
It is easy to deduce that for any $u\in\Lambda$, we have $u_{\lambda}\in\Lambda$ for all $\lambda>0$. Consequently, $\inf_{\lambda>0}I(u_{\lambda})\leq I(u_{1})=I(u)$, which implies
(4.4) $\inf_{u\in\Lambda}\inf_{\lambda>0}I(u_{\lambda})\leq\inf_{u\in\Lambda}I(u).$
Now set $a=\alpha p-(d-sp)>0$, $b=d-\alpha q>0$, $M=\frac{1}{p}[u]^{p}_{W^{s,p}(\mathbb{R}^{d})}>0$, and $N=\frac{1}{q}\|u\|^{q}_{L^{q}(\mathbb{R}^{d})}>0$. By Lemma 4.1, there exists a unique $\lambda_{u}>0$ such that $\inf_{\lambda>0}I(u_{\lambda})=I(u_{\lambda_{u}})$ for every $u\in\Lambda$. Therefore, using the fact that for any $u\in\Lambda$ we have $u_{\lambda_{u}}\in\Lambda$, we obtain
$\inf_{u\in\Lambda}I(u)\leq I(u_{\lambda_{u}})=\inf_{\lambda>0}I(u_{\lambda}),\quad\forall\,u\in\Lambda\quad\Rightarrow\quad\inf_{u\in\Lambda}I(u)\leq\inf_{u\in\Lambda}\inf_{\lambda>0}I(u_{\lambda}).$
Applying Lemma 4.1 once again with the above parameters, and using the two estimates obtained above, we conclude that
$\inf_{u\in\Lambda}I(u)$ $=\inf_{u\in\Lambda}\,\inf_{\lambda>0}I(u_{\lambda})$
$=\inf_{u\in\Lambda}\,\inf_{\lambda>0}\left\{\lambda^{\alpha p-(d-sp)}\frac{1}{p}[u]^{p}_{W^{s,p}(\mathbb{R}^{d})}+\lambda^{-(d-\alpha q)}\frac{1}{q}\|u\|^{q}_{L^{q}(\mathbb{R}^{d})}\right\}$
$=\frac{\alpha(p-q)+sp}{\alpha p-(d-sp)}\left(\frac{d-\alpha q}{\alpha p-(d-sp)}\right)^{\frac{\alpha q-d}{\alpha(p-q)+sp}}\left(\frac{1}{p^{\frac{d-\alpha q}{\alpha(p-q)+sp}}}\right)\left(\frac{1}{q^{\frac{\alpha p-(d-sp)}{\alpha(p-q)+sp}}}\right)\times$
$\hskip 14.22636pt\inf_{u\in\Lambda}[u]^{{\frac{p(d-\alpha q)}{\alpha(p-q)+sp}}}_{W^{s,p}(\mathbb{R}^{d})}\,\|u\|^{\frac{q(\alpha p-(d-sp))}{\alpha(p-q)+sp}}_{L^{q}(\mathbb{R}^{d})}.$
Let
$\widetilde{C}:=\frac{\alpha(p-q)+sp}{\alpha p-(d-sp)}\left(\frac{d-\alpha q}{\alpha p-(d-sp)}\right)^{\frac{\alpha q-d}{\alpha(p-q)+sp}}\left(\frac{1}{p^{\frac{d-\alpha q}{\alpha(p-q)+sp}}}\right)\left(\frac{1}{q^{\frac{\alpha p-(d-sp)}{\alpha(p-q)+sp}}}\right).$
Then, we have
$\inf_{u\in\Lambda}I(u)=\widetilde{C}\,\inf_{u\in\Lambda}\,[u]^{{\frac{p(d-\alpha q)}{\alpha(p-q)+sp}}}_{W^{s,p}(\mathbb{R}^{d})}\,\|u\|^{\frac{q(\alpha p-(d-sp))}{\alpha(p-q)+sp}}_{L^{q}(\mathbb{R}^{d})}.$
For any $u\in\mathcal{D}^{p,q}_{s}$, define $u_{0}(x)=K^{\frac{1}{r}}\frac{u(x)}{\|u\|_{L^{r}(\mathbb{R}^{d})}}$. Then,
$\int_{\mathbb{R}^{d}}|u_{0}(x)|^{r}\,dx=K.$
Therefore, $u_{0}\in\Lambda$. Then, we have
$\inf_{u\in\Lambda}\,I(u)$ $=\widetilde{C}\,\inf_{u\in\mathcal{D}^{p,q}_{s}}\,\left[\frac{K^{\frac{1}{r}}u}{\|u\|_{L^{r}(\mathbb{R}^{d})}}\right]^{{\frac{p(d-\alpha q)}{\alpha(p-q)+sp}}}_{W^{s,p}(\mathbb{R}^{d})}\,\left\|\frac{K^{\frac{1}{r}}u}{\left\|u\|_{L^{r}(\mathbb{R}^{d})}\right\|}\right\|^{\frac{q(\alpha p-(d-sp))}{\alpha(p-q)+sp}}_{L^{q}(\mathbb{R}^{d})}$
$=\widetilde{C}\,K^{\frac{d(p-q)+spq}{r(\alpha(p-q)+sp)}}\,\inf_{u\in\mathcal{D}^{p,q}_{s}}\,\frac{[u]^{{\frac{p(d-\alpha q)}{\alpha(p-q)+sp}}}_{W^{s,p}(\mathbb{R}^{d})}\,\|u\|^{\frac{q(\alpha p-(d-sp))}{\alpha(p-q)+sp}}_{L^{q}(\mathbb{R}^{d})}}{\|u\|^{\frac{d(p-q)+spq}{\alpha(p-q)+sp}}_{L^{r}(\mathbb{R}^{d})}}.$
Next, we fix
$K=\left(\frac{1}{\widetilde{C}}\right)^{\frac{r(\alpha(p-q)+sp)}{d(p-q)+spq}}$ $=\left(\frac{\alpha p-(d-sp)}{\alpha(p-q)+sp}\right)^{\frac{r(\alpha(p-q)+sp)}{d(p-q)+spq}}\left(\frac{\alpha p-(d-sp)}{d-\alpha q}\right)^{\frac{r(\alpha q-d)}{d(p-q)+spq}}\times$
$\hskip 17.07164ptp^{\frac{r(d-\alpha q)}{d(p-q)+spq}}\,q^{\frac{r(\alpha p-(d-sp))}{d(p-q)+spq}},$
which implies that
$\eta=\inf_{u\in\Lambda}\,I(u)=\inf_{u\in\mathcal{D}^{p,q}_{s}}\,\frac{[u]^{{\frac{p(d-\alpha q)}{\alpha(p-q)+sp}}}_{W^{s,p}(\mathbb{R}^{d})}\,\|u\|^{\frac{q(\alpha p-(d-sp))}{\alpha(p-q)+sp}}_{L^{q}(\mathbb{R}^{d})}}{\|u\|^{\frac{d(p-q)+spq}{\alpha(p-q)+sp}}_{L^{r}(\mathbb{R}^{d})}}.$
Therefore, we obtain
$\|u\|_{L^{r}(\mathbb{R}^{d})}\leq\left(\frac{1}{\eta}\right)^{\frac{\alpha(p-q)+sp}{d(p-q)+spq}}\,[u]^{\frac{p(d-\alpha q)}{d(p-q)+spq}}_{W^{s,p}(\mathbb{R}^{d})}\,\|u\|^{\frac{q(\alpha p-(d-sp))}{d(p-q)+spq}}_{L^{q}(\mathbb{R}^{d})},\quad\forall\ u\in\mathcal{D}^{p,q}_{s}.$
Define $\mathscr{L}:=\left(\frac{1}{\eta}\right)^{\frac{\alpha(p-q)+sp}{d(p-q)+spq}}$. From the above analysis, it follows that, $\mathscr{L}$ is the optimal constant in the interpolation inequality. Substituting the value of $\alpha$ in the above inequality, we arrive at
$\|u\|_{L^{r}(\mathbb{R}^{d})}\leq\mathscr{L}[u]^{a}_{W^{s,p}(\mathbb{R}^{d})}\|u\|^{1-a}_{L^{q}(\mathbb{R}^{d})},\quad\forall\ u\in\mathcal{D}^{p,q}_{s}.$
This proves the theorem. ∎
Acknowledgement: The author thanks the Department of Mathematics and Statistics at the Indian Institute of Technology Kanpur, India, for providing a supportive research environment. The author also acknowledges the support of the Indian Institute of Technology Kanpur, India, through the FARE (Fellowship for Academic and Research Excellence) fellowship. The author is thankful to Michał Kijaczko for suggesting the extension of logarithmic Sobolev inequalities to the weighted fractional Sobolev space.